### v1

@@ Abstract
Explainability is a key component in many applications involving deep neural networks (DNNs). However, current explanation methods for DNNs commonly leave it to the human observer to distinguish relevant explanations from spurious noise. This is not feasible anymore when going from easily human-accessible data such as images to more complex data such as genome sequences. To facilitate the accessibility of DNN outputs from such complex data and to increase explainability, we present a modification of the widely used explanation method layer-wise relevance propagation. Our approach enforces sparsity directly by pruning the relevance propagation for the different layers. Thereby, we achieve sparser relevance attributions for the input features as well as for the intermediate layers. As the relevance propagation is input-specific, we aim to prune the relevance propagation rather than the underlying model architecture. This allows to prune different neurons for different inputs and hence, might be more appropriate to the local nature of explanation methods. To demonstrate the efficacy of our method, we evaluate it on two types of data, images and genomic sequences. We show that our modification indeed leads to noise reduction and concentrates relevance on the most important features compared to the baseline.
@@
We gratefully acknowledge funding by grants KL 3037/7-1 (to NK) and RE 3474/8-1 (to BYR), project P5 in the Research Unit KI-FOR 5363 of the German Research Foundation (DFG).
@@ 1 Introduction
As the usage of deep neural networks (DNNs) has grown tremendously in a variety of fields, so has the interest in the explainability and interpretability of such models. Especially for sensitive areas such as medical imaging (Litjens et al.,, 2017) or genomics (Eraslan et al.,, 2019, Bartoszewicz et al.,, 2021), an understanding of how the model came to certain predictions is essential to build and ensure trustworthiness. Therefore, so-called post-hoc attribution or explanation methods have been introduced (Samek et al.,, 2021) to decode the black-box behavior common for DNN architectures. Thereby, we assume there is an already trained model and consider a separate method to explain it. Instead of global explanations of the model, our focus is on local methods. Their general idea is to obtain an input-specific explanation of the decisive behavior of the model by attributing relevance scores to every input dimension based on the model’s prediction. It has been shown that several of these methods are special cases of the more general game-theoretic concept of Shapley values (Lundberg and Lee,, 2017). Layer-wise relevance propagation (LRP; Bach et al.,, 2015) is among these methods and propagates relevance backward through the network from one layer to another. The basic LRP approach was subsequently developed and extended to further network architectures (Binder et al.,, 2016, Arras et al.,, 2017, Schnake et al.,, 2021, Ali et al.,, 2022, Montavon et al.,, 2019). It has gained increased popularity since then.
@@
For image classification, a human evaluation and interpretation of the LRP explanation can be done qualitatively by a suitable visualization such as heat-maps. The observer can identify the object to be detected and decide whether the corresponding relevance attribution split across the input dimensions matches this object. However, for high-dimensional data with unknown ground truth, this might not be feasible anymore. For example, this is often the case for genome sequences. Besides the high-dimensionality, noisy relevance attributions can make it difficult to identify and interpret the main drivers for the model’s prediction in such data and hence, diminish the meaningfulness of the explanation. As a solution, we propose a modification of LRP that generates sparser explanations by extending the idea of filtering or pruning the relevance propagation (Montavon et al.,, 2018). We call our approach pruned layer-wise relevance propagation (PLRP). Sparsification of the explanation might be desirable in the sense that it reduces noise and the number of features with non-zero relevance, i.e., highlights only the most important features. Furthermore, our method is able to regulate the degree of sparsity, which can be beneficial for feature selection or the identification of concepts learned by the network in deeper layers. The basic intuition of our method is illustrated in Figure 1(a) via three heatmaps of images obtained through LRP (left) and their corresponding sparsified versions through PLRP (right).
Further, Figure 1(b) shows how the relevance propagation in our proposed method PLRP is pruned within a DNN by removing less relevant neurons. This idea of maintaining only the most important features is inspired by $L1$-regularization. However, integrating $L1$-regularization into LRP is not directly applicable. This is because the map representing the relevance propagation from one layer to another, defined by LRP, is input-specific. Hence, there is no single consistent map for all data points. Approximating all inputs with the same model would be a strong simplification of the original explanation method. It would also disregard the fact that different inputs might activate different neurons in intermediate layers. Altogether, this motivates our contribution of PLRP.
@@ 2 Related Work
Although related to model pruning (Zhu and Gupta,, 2017), our approach and goal of pruning relevance propagation conceptually differ from it. Instead of pruning the underlying model, we aim to sparsify the individual explanations, which are local approximations of the model. Hence, we prune different neurons depending on the input features, thereby enabling a local and input-specific pruning. In contrast, when Gupta et al., (2021) and Yeom et al., (2021) applied LRP for model pruning, they used the relevance attribution aggregated over multiple inputs to prune the model globally. The authors show that the number of parameters can be reduced drastically, while still maintaining or even improving the predictive performance of the model.
@@ 3.1 Setting
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C},\boldsymbol{x}=(x_{1},\ldots,x_{d})^{\top}\mapsto f(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),\ldots,f_{C}(\boldsymbol{x}))^{\top}$ be the pre-softmax prediction function of a trained DNN for a classification task with $C$ classes, which we also refer to as model. For our setting, we are only interested in the prediction score of the winning class, denoted by $f_{c^{*}}(\boldsymbol{x})$. For a given model, an explanation method is a function $e:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}:\boldsymbol{x}\mapsto e(\boldsymbol{x})=(r_{1},...,r_{d})^{\top}$ that assigns a relevance score $r_{i}$ to every input dimension $i=1,\ldots,d$. We denote by $W_{l}\in\mathbb{R}^{m\times n}$ the weight matrix of the $l$-th layer of the model and by $w_{jk}^{(l)}$ the weight of the edge connecting the $j$-th neuron in layer $l-1$ with the $k$-th neuron in layer $l$. The corresponding activations are denoted by $a_{j}^{(l-1)}$ and $a_{k}^{(l)}$, respectively. Hence, we can write $\boldsymbol{a}^{(l)}=\big{(}W^{\top}_{l}\boldsymbol{a}^{(l-1)}\big{)}^{+}$, where $(\cdot)^{+}$ refers to the entry-wise application of the ReLU function. Since for our purposes we consider only one layer at a time, we will drop the layer-specific superscripts for readability and simply write $w_{jk}$. Furthermore, let $(\cdot)^{-}$ be the function that only keeps the absolute value of the negative part and is zero otherwise.
@@
As described earlier, LRP propagates the relevance backward from one layer to its predecessor, starting with the prediction score of the output layer $f_{c^{*}}(\boldsymbol{x})$. The total relevance mass is preserved in every layer by the so-called conservation property (Bach et al.,, 2015). The relevance propagation from layer $l$ to $l-1$ can be expressed by a linear map (Sixt et al.,, 2020), i.e., a multiplication with the matrix $M_{l}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}:\boldsymbol{r}^{(l)}\mapsto\boldsymbol{r}^{(l-1)}=M_{l}\boldsymbol{r}^{(l)}$. Thereby, the matrix entry in the $i$-th row and $j$-th column represents the proportion of relevance of neuron $j$ in layer $l$ that is allocated to neuron $i$ in layer $l-1$. As for the weights, we will disregard the layer subscript and only write $M$. There are different LRP rules for how these proportions are calculated (Montavon et al.,, 2018). Specifically, for the LRP-0 rule, the matrix $M$ is given by
@@ 3.2 Pruned LRP
The goal of pruned layer-wise relevance propagation (PLRP) is to create sparser explanations by reducing noise and simultaneously increasing the attribution to the most relevant features. We do this by pruning the explanation directly in every layer. This consists of two steps: first, determining the neurons to be pruned, and second, redistributing their relevance mass among the remaining neurons. The criterion for pruning the neurons is the relevance itself, i.e., the neurons with the lowest relevance scores get pruned. Hence, first, we conduct a preliminary unpruned relevance attribution to layer $l$, prune it, and then after redistribution, use it as input for the attribution to layer $l-1$. As it is unclear in advance how the relevance is distributed among the neurons, we prune a proportion of the total relevance mass instead of a proportion of the neurons.
@@ 3.2.1 Pruning Relevance
We denote the unpruned relevance vector by $\widetilde{\boldsymbol{r}}^{(l-1)}$. Then for a pre-defined proportion $p_{l}\in[0,1]$ of the total relevance mass that is supposed to be pruned, we determine the corresponding threshold $\theta_{l}\in\mathbb{R}$, such that the relevance scores below $\theta_{l}$ sum up to $p_{l}$ of the total relevance mass. Hence, for (unpruned) ordered relevance scores $\widetilde{r}_{s_{1}}^{(l-1)}\leq\widetilde{r}_{s_{2}}^{(l-1)}\leq\ldots\leq\widetilde{r}_{s_{n}}^{(l-1)}$, we set
@@ 3.2.2 Redistributing Pruned Relevance
[…]
@@
Hence, by attributing relevance according to $\widehat{M}_{l}$, we ensure that neurons receive zero relevance that would have been assigned a relevance mass below the threshold $\theta_{l}$ of Equation 3.2 otherwise. As $\widehat{M}_{l}$ is also a matrix in the form of Equation 3.1, only with more zero activations, the pruned relevance gets implicitly redistributed in the manner of LRP.
@@ 4 Evaluation
We evaluate the two variants of PLRP in the context of classification of two different data types: images and genomic sequences. For the quantitative evaluation, we use multiple metrics covering different aspects of the explanation method following the categorization of Hedström et al., (2023). Further, we evaluate our methods qualitatively by considering heatmaps and sequence logo plots of the relevance attributions.
@@ 4.1 Evaluation Metrics
Sparsity and localization are intended to measure the efficacy of our method, while faithfulness and robustness serve as sanity checks for its trustworthiness.
@@
As we are interested in sparse explanations that concentrate the relevance mass at few important features, we evaluate the degree of sparsity or complexity by the Gini Index and entropy of the relevance attribution as proposed by Chalasani et al., (2020) and Bhatt et al., (2020). Thereby, a higher Gini Index or lower entropy, respectively, indicate that there are more zero entries and more mass is concentrated at fewer entries.
@@
Localization measures how accurately the relevance is attributed to a pre-specified region. Hence, for this metric, a known ground truth is needed. We therefore calculate the Relevance Mass Accuracy (RMA; Arras et al.,, 2022), which measures how much relevance mass out of the total mass is allocated to the ground truth.
@@
Faithfulness measures whether the explanation method indeed uncovers the predictive behavior of the underlying model. This means that the highest relevance scores correspond to the most decisive input features. The idea is to perturb the input based on the relevance attribution and then measure how quickly the predictive score drops. This procedure is also referred to as pixel-flipping (Bach et al.,, 2015) or selectivity (Samek et al.,, 2016). Hence, we have a perturbation function that takes the model input $\boldsymbol{x}$ and corresponding relevance attribution $\boldsymbol{r}=e(\boldsymbol{x})$ and outputs a sequence of perturbed versions of $\boldsymbol{x}$ based on $\boldsymbol{r}$, i.e., $(\boldsymbol{x},\boldsymbol{r})\mapsto(\boldsymbol{x}_{1}^{\prime},\ldots,\boldsymbol{x}_{s}^{\prime})$ where $s$ denotes the number of perturbation steps. As the actual value of prediction scores can differ from one class to another, we normalize the scores and measure the relative drop. Further, to compare the results across different pruning parameters $p_{l}$, we also consider the area under the curve (AUC). Note that a quicker decrease in the prediction score corresponds to a lower AUC and hence, to a better explanation with respect to this metric.
Robustness
For explanation methods, robustness refers to the idea that similar inputs should produce similar explanations. Hence, it can be seen as a sensitivity proxy that measures how much the explanation changes for a small shift in the input. Alvarez-Melis and Jaakkola, (2018) propose to estimate the local Lipschitz constant, i.e., for $\varepsilon>0$ and some input $\boldsymbol{x}\in\mathbb{R}^{d}$, the aim is to find a constant $L=L(\boldsymbol{x})$ such that for all $\boldsymbol{x}^{\prime}\in B_{\varepsilon}(\boldsymbol{x})=\{\boldsymbol{x}^{\prime}:\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|<\varepsilon\}$ it holds $\|e(\boldsymbol{x})-e(\boldsymbol{x}^{\prime})\|<L\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|$. Hence, $L$ is the steepest possible ascent of $e$ in a small neighborhood of $\boldsymbol{x}$. As calculating the gradient of the explanation method might be very expensive, or it might not even be differentiable, $L$ is estimated numerically by sampling from $B_{\varepsilon}(\boldsymbol{x})$.

### v2

@@ Abstract
Explainability is a key component in many applications involving deep neural networks (DNNs). However, current explanation methods for DNNs commonly leave it to the human observer to distinguish relevant explanations from spurious noise. This is not feasible anymore when going from easily human-accessible data such as images to more complex data such as genome sequences. To facilitate the accessibility of DNN outputs from such complex data and to increase explainability, we present a modification of the widely used explanation method layer-wise relevance propagation. Our approach enforces sparsity directly by pruning the relevance propagation for the different layers. Thereby, we achieve sparser relevance attributions for the input features as well as for the intermediate layers. As the relevance propagation is input-specific, we aim to prune the relevance propagation rather than the underlying model architecture. This allows to prune different neurons for different inputs and hence, might be more appropriate to the local nature of explanation methods. To demonstrate the efficacy of our method, we evaluate it on two types of data: images and genome sequences. We show that our modification indeed leads to noise reduction and concentrates relevance on the most important features compared to the baseline.
@@
We gratefully acknowledge funding by grants KL 3037/7-1 (to NK) and RE 3474/8-1 (to BYR), project P5 in the Research Unit KI-FOR 5363 (grant 459422098) of the German Research Foundation (DFG). First published in Machine Learning and Knowledge Discovery in Databases. Research Track. ECML PKDD 2024. Lecture Notes in Computer Science, Vol. 14944, pp 336–351, 2024 by Springer Nature.
@@ 1 Introduction
As the usage of deep neural networks (DNNs) has grown tremendously in a variety of fields, so has the interest in the explainability and interpretability of such models. Especially for sensitive areas such as medical imaging (Litjens et al.,, 2017) or genomics (Eraslan et al.,, 2019, Bartoszewicz et al.,, 2021), an understanding of how the model came to certain predictions is essential to build and ensure trustworthiness. Therefore, so-called post-hoc attribution or explanation methods have been introduced (Samek et al.,, 2021) to decode the black-box behavior common for DNN architectures. Thereby, we assume there is an already trained model and consider a separate method to explain it. Instead of global explanations of the model, our focus is on local methods. Their general idea is to obtain an input-specific explanation of the decisive behavior of the model by attributing relevance scores to every input dimension based on the model’s prediction. It has been shown that several of these methods are special cases of the more general game-theoretic concept of Shapley values (Lundberg and Lee,, 2017). Layer-wise relevance propagation (LRP; Bach et al.,, 2015) is among these methods and propagates relevance backward through the network from one layer to another. The basic LRP approach was subsequently developed and extended to further network architectures (Binder et al.,, 2016, Arras et al.,, 2017, Schnake et al.,, 2021, Ali et al.,, 2022, Montavon et al.,, 2019). It has gained increased popularity since then and is widely used to explain a model’s prediction.
@@
For image classification, a human evaluation and interpretation of the LRP explanation can be done qualitatively by a suitable visualization such as heatmaps. The observer can identify the object to be detected and decide whether the corresponding relevance attribution split across the input dimensions matches this object. However, for high-dimensional data with unknown ground truth, this might not be feasible anymore. For example, this is often the case for genome sequences. Besides the high dimensionality, noisy relevance attributions can make it difficult to identify and interpret the main drivers for the model’s prediction in such data and hence, diminish the meaningfulness of the explanation. As a solution, we propose a modification of LRP that generates sparser explanations by extending the idea of filtering or pruning the relevance propagation (Montavon et al.,, 2018). We call our approach pruned layer-wise relevance propagation (PLRP). Sparsification of the explanation might be desirable in the sense that it reduces noise and the number of features with non-zero relevance, i.e., highlights only the most important features. Furthermore, our method is able to regulate the degree of sparsity, which can be beneficial for feature selection or the identification of concepts learned by the network in deeper layers. While our approach is specific to LRP, the general idea of enforcing sparsity is not. In fact, a similar modification might be applicable to other local explanation methods. The basic intuition of our method is illustrated in Figure 1(a) via three heatmaps of images obtained through LRP (left) and their corresponding sparsified versions through PLRP (right). Further, Figure 1(b) shows how the relevance propagation in our proposed method PLRP is pruned within a DNN by removing less relevant neurons. This idea of maintaining only the most important features is inspired by $L1$-regularization. However, integrating $L1$-regularization into LRP is not directly applicable. This is because the map representing the relevance propagation from one layer to another, defined by LRP, is input-specific. Hence, there is no single consistent map for all data points. Approximating all inputs with the same model would be a strong simplification of the original explanation method. It would also disregard the fact that different inputs might activate different neurons in intermediate layers. Altogether, this motivates our contribution of PLRP.
@@ 2 Related Work
Adebayo et al., (2018) and Sixt et al., (2020) conclude that some explanation methods, such as LRP, might be insensitive to changes in the model’s parameters and hence, question their reliability. However, in a later study, Arras et al., (2022) cannot confirm these results. In fact, in their evaluation, LRP is among the best-performing explanation methods. The authors argue the insensitivity might be more likely driven by non-class-discriminative explanations or only asymptotically relevant.
Although related to model pruning (Zhu and Gupta,, 2017), our approach and goal of pruning relevance propagation differ conceptually from it. Instead of pruning the underlying model, we aim to sparsify the individual explanations, which are local approximations of the model. Hence, we prune different neurons depending on the input features, thereby enabling a local and input-specific pruning. In contrast, when Gupta et al., (2021) and Yeom et al., (2021) applied LRP for model pruning, they used the relevance attribution aggregated over multiple inputs to prune the model globally. The authors show that the number of parameters can be reduced drastically, while still maintaining or even improving the predictive performance of the model.
@@ 3.1 Setting
Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C},\boldsymbol{x}=(x_{1},\ldots,x_{d})^{\top}\mapsto f(\boldsymbol{x})=(f_{1}(\boldsymbol{x}),\ldots,f_{C}(\boldsymbol{x}))^{\top}$ be the pre-softmax prediction function of a trained DNN for a classification task with $C$ classes, which we also refer to as model. For our setting, we are only interested in the prediction score of the winning class, denoted by $f_{c^{*}}(\boldsymbol{x})$. For a given model, an explanation method is a function $e:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}:\boldsymbol{x}\mapsto e(\boldsymbol{x})=(r_{1},...,r_{d})^{\top}$ that assigns a relevance score $r_{i}$ to every input dimension $i=1,\ldots,d$. We denote by $W_{l}\in\mathbb{R}^{m\times n}$ the weight matrix of the $l$-th layer of the model and by $w_{jk}^{(l)}$ the weight of the edge connecting the $j$-th neuron in layer l1l-1 with the $k$-th neuron in layer $l$. The corresponding activations are denoted by $a_{j}^{(l-1)}$ and $a_{k}^{(l)}$, respectively. Hence, we can write $\boldsymbol{a}^{(l)}=\big(W^{\top}_{l}\boldsymbol{a}^{(l-1)}\big)^{+}$, where $(\cdot)^{+}$ refers to the entry-wise application of the ReLU function. Since for our purposes we consider only one layer at a time, we will drop the layer-specific superscripts for readability and simply write $w_{jk}$. Furthermore, let $(\cdot)^{-}$ be the function that only keeps the absolute value of the negative part and is zero otherwise.
@@
As described earlier, LRP propagates the relevance backward from one layer to its predecessor, starting with the prediction score of the output layer $f_{c^{*}}(\boldsymbol{x})$. The total relevance mass is preserved in every layer by the so-called conservation property (Bach et al.,, 2015). The relevance propagation from layer $l$ to l1l-1 can be expressed by a linear map (Sixt et al.,, 2020), i.e., a multiplication with the matrix $M_{l}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}:\boldsymbol{r}^{(l)}\mapsto\boldsymbol{r}^{(l-1)}=M_{l}\boldsymbol{r}^{(l)}$. Thereby, the matrix entry in the $i$-th row and $j$-th column represents the proportion of relevance of neuron $j$ in layer $l$ that is allocated to neuron $i$ in layer l1l-1. As for the weights, we will disregard the layer subscript and only write $M$. There are different LRP rules for how these proportions are calculated (Montavon et al.,, 2018). Specifically, for the LRP-0 rule, the matrix $M$ is given by
@@ 3.2 Pruned LRP
The goal of pruned layer-wise relevance propagation (PLRP) is to create sparser explanations by reducing noise and simultaneously increasing the attribution to the most relevant features. We do this by pruning the explanation directly in every layer. This consists of two steps: first, determining the neurons to be pruned, and second, redistributing their relevance mass among the remaining neurons. The criterion for pruning the neurons is the relevance itself, i.e., the neurons with the lowest relevance scores get pruned. Hence, first, we conduct a preliminary unpruned relevance attribution to layer $l$, prune it, and then after redistribution, use it as input for the attribution to layer l1l-1. As it is unclear in advance how the relevance is distributed among the neurons, we prune a proportion of the total relevance mass instead of a proportion of the neurons.
@@ 3.2.1 Pruning Relevance
We denote the unpruned relevance vector by $\widetilde{\boldsymbol{r}}^{(l-1)}$. Then for a pre-defined proportion $p_{l}\in[0,1)$ of the total relevance mass that is supposed to be pruned, we determine the corresponding threshold $\theta_{l}\in\mathbb{R}$, such that the relevance scores below $\theta_{l}$ sum up to $p_{l}$ of the total relevance mass. Hence, for (unpruned) ordered relevance scores $\widetilde{r}_{s_{1}}^{(l-1)}\leq\widetilde{r}_{s_{2}}^{(l-1)}\leq\ldots\leq\widetilde{r}_{s_{n}}^{(l-1)}$, we set
@@ 3.2.2 Redistributing Pruned Relevance
$\boldsymbol{r}^{(l-1)}:=\lambda_{l}\,\boldsymbol{1}_{\big\{\widetilde{\boldsymbol{r}}^{(l-1)}>\theta_{l}\big\}}\widetilde{\boldsymbol{r}}^{(l-1)}\,$,
@@
Hence, by attributing relevance according to $\widehat{M}_{l}$, we ensure that neurons receive zero relevance that would have been assigned a relevance mass below the threshold $\theta_{l}$ of Equation (3.2) otherwise. As $\widehat{M}_{l}$ is also a matrix in the form of Equation (3.1), only with more zero activations, the pruned relevance gets implicitly redistributed in the manner of LRP.
@@ 4 Evaluation
We evaluate the two variants of PLRP in the context of classification of two different data types: images and genome sequences. For the quantitative evaluation, we use multiple metrics covering different aspects of the explanation method following the categorization of Hedström et al., (2023). Further, we evaluate our methods qualitatively by considering heatmaps and sequence logo plots of the relevance attributions.
@@ 4.1 Evaluation Metrics
Sparsity and localization are intended to measure the efficacy of our method, while robustness and faithfulness serve as sanity checks for its trustworthiness. Note that since these metrics consider different aspects, they do not necessarily align. $\uparrow$ and $\downarrow$ hint whether a higher or lower value indicate better outcomes.
@@
As we are interested in sparse explanations that concentrate the relevance mass at few important features, we evaluate the degree of sparsity or complexity by the Gini Index ($\uparrow$) and entropy ($\downarrow$) of the relevance attribution as proposed by Chalasani et al., (2020) and Bhatt et al., (2020). Thereby, a higher Gini Index or lower entropy, respectively, indicate that there are more zero entries and more mass is concentrated at fewer entries.
@@
Localization measures how accurately the relevance is attributed to a pre-specified region. Hence, for this metric, a known ground truth is needed. Therefore, we calculate the Relevance Mass Accuracy (RMA, $\uparrow$; Arras et al.,, 2022), which measures how much relevance mass out of the total mass is allocated to the ground truth.
Robustness
For explanation methods, robustness refers to the idea that similar inputs should produce similar explanations. Hence, it can be seen as a sensitivity proxy that measures how much the explanation changes for a small shift in the input. Alvarez-Melis and Jaakkola, (2018) propose to estimate the local Lipschitz constant ($\downarrow$), i.e., for $\varepsilon>0$ and some input $\boldsymbol{x}\in\mathbb{R}^{d}$, the aim is to find a constant $L=L(\boldsymbol{x})$ such that for all $\boldsymbol{x}^{\prime}\in B_{\varepsilon}(\boldsymbol{x})=\{\boldsymbol{x}^{\prime}:\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|<\varepsilon\}$ it holds $\|e(\boldsymbol{x})-e(\boldsymbol{x}^{\prime})\|<L\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|$. Hence, $L$ is the steepest possible ascent of $e$ in a small neighborhood of $\boldsymbol{x}$. As calculating the gradient of the explanation method might be very expensive, or $e$ might not even be differentiable, $L$ is estimated numerically by sampling from $B_{\varepsilon}(\boldsymbol{x})$.
@@
Faithfulness measures whether the explanation method indeed uncovers the predictive behavior of the underlying model. This means that the highest relevance scores correspond to the most decisive input features. The idea is to perturb the input based on the relevance attribution and then measure how quickly the predictive score drops. This procedure is also referred to as pixel-flipping (Bach et al.,, 2015) or selectivity (Samek et al.,, 2016). Hence, we have a perturbation function that takes the model input $\boldsymbol{x}$ and corresponding relevance attribution $\boldsymbol{r}=e(\boldsymbol{x})$ and outputs a sequence of perturbed versions of $\boldsymbol{x}$ based on $\boldsymbol{r}$, i.e., $(\boldsymbol{x},\boldsymbol{r})\mapsto(\boldsymbol{x}_{1}^{\prime},\ldots,\boldsymbol{x}_{s}^{\prime})$ where $s$ denotes the number of perturbation steps. As the actual value of prediction scores can differ from one class to another, we normalize the scores and measure the relative drop. Further, to compare the results across different pruning parameters $p_{l}$, we also consider the area under the curve (AUC, $\downarrow$). Note that a quicker decrease in the prediction score corresponds to a lower AUC and hence, to a better explanation with respect to this metric. As faithfulness might be sensitive to the choice of perturbation, e.g., size of perturbed area or type of noise to be added, this metric should be considered cautiously. This holds especially when comparing small differences in this metric. Nevertheless, it can serve as a proxy for the general trend in how faithful a method is.